# Resolving Gauge Ambiguities of the Berry Connection in Non-Hermitian Systems

### v1

@@ VI Conclusions
In this work, we have introduced a unique CBC that guarantees the norm preservation of quantum states governed by non-Hermitian Hamiltonians. By incorporating the parameter-dependent Hilbert-space metric and associated with it the covariant derivative, the resulting covariant framework properly accounts for the general ${\rm GL}(N,\mathbb{C})$ frame transformations, ensuring that the CBC remains real-valued, and eliminates gauge ambiguities inherent to conventional biorthogonal geometric constructions. This provides a consistent and unambiguous geometric foundation for defining Berry phases, non-Abelian holonomies, and topological invariants in genuinely quantum non-Hermitian systems.
We additionally illustrated the proposed framework using a two-level pseudo-Hermitian model describing topological magnons on a honeycomb lattice [8], demonstrating that the associated Berry curvature, stemming from the covariant Berry connection, is everywhere flat in the unbroken pseudo-Hermitian phase. From this perspective, previously reported emergent (2+1)-dimensional electromagnetic fields stemming from nonzero Berry curvature in such systems [9] remain relevant in (semi)classical regime, where the requirement of quantum-state norm preservation can be relaxed.

### v2

@@ VI Conclusions
In this work, we have introduced a uniquely defined, real-valued Berry connection that characterizes the geometric and topological structure of the eigenspace of non-Hermitian Hamiltonians in the genuine quantum regime. By incorporating the parameter-dependent Hilbert-space metric and associated with it the covariant derivative, the resulting covariant framework properly accounts for the general ${\rm GL}(N,\mathbb{C})$ frame transformations, ensuring that the CBC remains real-valued, and eliminates gauge ambiguities inherent to the conventional biorthogonal Berry connection. This provides a consistent and unambiguous geometric framework for defining both the Abelian and non-Abelian holonomies, curvatures, and topological invariants in quantum systems described by non-Hermitian Hamiltonians. We additionally illustrated the proposed framework using a known two-level pseudo-Hermitian model, demonstrating that the geometric structure encoded in its Hamiltonian eigenspace exhibits a substantial qualitative difference between the classical and quantum regimes. This distinction highlights the central role of quantum-state norm preservation, properly incorporated via the covariant formalism, in defining a physically consistent Berry connection and curvature for non-Hermitian systems.